Suppose that a population of women heights is normally distributed with mean of 68



inches and standard deviation of 3 inches. If a person is selected at random, what is the



probability that her height is less than 66 inches?

1
Expert's answer
2022-05-04T01:43:20-0400

Denote by "X" a random variable, which is normally distributed with parameters "\mu=68" and "\sigma=3". The probability density function is: "p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}=\frac{1}{3\sqrt{2\pi}}e^{-\frac12\left(\frac{x-68}{3}\right)^2}". The probability is: "P(X\leq66)=\int_{-\infty}^{66}\frac{1}{3\sqrt{2\pi}}e^{-\frac12\left(\frac{x-68}{3}\right)^2}dx\approx0.2525". The answer is rounded to 4 decimal places.

Answer: "0.2525" (rounded to 4 decimal places).